OpenStudy (nikki_skellington):

If (7^2)^x = 1, what is the value of x? Explain your answer. If (7^0)^x = 1, what are the possible values of x? Explain your answer.

OpenStudy (vuriffy):

What's 7*7?

OpenStudy (nikki_skellington):

49

OpenStudy (vuriffy):

You may rewrite it as (49)^x = 1.

OpenStudy (nikki_skellington):

w8 is that for the first one

OpenStudy (vuriffy):

Yes.

OpenStudy (nikki_skellington):

shouldn't it be 7 times 2 or no?

OpenStudy (vuriffy):

Do you know what exponent always makes the numerical value as 1 for the result?

OpenStudy (vuriffy):

You wrote 7^2?

OpenStudy (vuriffy):

That's 7 to the power of 2. Which means (7)*(7) = 49.

OpenStudy (nikki_skellington):

ohhh omg im so dumb -.-

OpenStudy (nikki_skellington):

ok so the answer would be 1 for the first correct

OpenStudy (vuriffy):

Take a step back.

OpenStudy (vuriffy):

If the exponent is 0, the value becomes 1.

OpenStudy (vuriffy):

If x was 1, it would be 49^1 = 49.

OpenStudy (vuriffy):

49 does not equal 1, sorry.

OpenStudy (nikki_skellington):

so w8 sorry im not good at this, the answer would be....49^1 = 49 or (49)^x = 1

OpenStudy (vuriffy):

Here, let me rewrite it: (7^2)^x = 1 Your GOAL is to find what the x value is which is an exponent in this case (49)^x = 1 If x = 1 (49)^1 = 1 49 does not = 1. if x = 0 (49)^0 = 1 1 = 1 Anything to the exponent of 1 equals 1.

OpenStudy (vuriffy):

Anything to the exponent of 0 equals 1, sorry for the typo.

OpenStudy (nikki_skellington):

ok i get it now thx! could u also help me with the second one?

OpenStudy (vuriffy):

Yes, of course, this one is actually cute, if I must say so myself.

OpenStudy (nikki_skellington):

lol ur a lifesaver i swear

OpenStudy (vuriffy):

(7^0)^x = 1 They try to trick you in a sense.

OpenStudy (vuriffy):

(7^0)^x = 1 7^0 already equals 1. (1)^x = 1. There is two values which allow this to equal one, can you guess the two values, or do you need me to explain it further?

OpenStudy (nikki_skellington):

sorry i dont know :/ would it be one of them be 0?

OpenStudy (vuriffy):

No, don't feel discouraged, you're fine. Yes, one would be 0, good job. (1)^0 = 1. 1 = 1. (:

OpenStudy (vuriffy):

Anything to the first power is the same numerical value. (1)^1 = 1 1 = 1 1*1 = 1. Just like when I stated (49)^1 = 49 49 * 1 = 49.

OpenStudy (nikki_skellington):

ohh ok so it would be 0 and 1 right

OpenStudy (vuriffy):

Yes, correct.

OpenStudy (nikki_skellington):

thank you soo much!

OpenStudy (vuriffy):

Anytime!